Solve by completing the square and separate answers with commas:
x^2 - 2x = 10
Please help Thanks!

Answer by stanbon(75887)

Solve by completing the square and separate answers with commas:
x^2 - 2x = 10
---------------------
x^2 - 2x + 1 = 10+1
(x-1)^2 = 11
x-1 = sqrt(11) or x-1 = -sqrt(11)
x = 1+sqrt(11) or x = 1-sqrt(11)
===================================
Cheers,
Stan H.
==========